OpenStudy (anonymous):

Height of Ball when punted = h=-0.1d^ +1.18d +2 a) what is the max height of this punt? b) how far will it have gone when it hits the ground?

OpenStudy (anonymous):

sorry equation is h=-.01d^ +1.18 +2

OpenStudy (anonymous):

still doesn't make sense... do you mean h=-.01d^2+1.18d+2?

OpenStudy (anonymous):

yes

myininaya (myininaya):

find derivative to find the max

OpenStudy (anonymous):

You find the max (which is the highest point which is called the vertex by) x = -b/(2a) x = -1.18/(2*-.01) x = 59

myininaya (myininaya):

or that

OpenStudy (anonymous):

Then you plug in x = 59 to find the height. y = -.01(59)^2 + 1.18(59) +2 y = -34.81 + 69.62 + 2 y = 36.81 Which is the height of the ball at the maximum point

OpenStudy (anonymous):

Thanks, Very Helpful!